OpenStudy (anonymous):

7. What is the surface area of a conical grain storage tank that has a height of 24 meters and a diameter of 16 meters? round the answer to the nearest square meter. 837 square meters 848 square meters 1,991 square meters 1,923 square meters

OpenStudy (anonymous):

Surface area: \(\Large S_a = \pi r h_s + \pi r^2 =\pi r \sqrt{r^2+h^2} + \pi r^2\) \(\Large S_a=\pi \times \frac{16}{2} \times \sqrt{(\frac{16}{2})^2 + 24^2} + \pi \times (\frac{16}{2})^2\) \(\Large S_a \approx 837\) \( \large m^2\)

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

@junghyunran is right!